OpenStudy (anonymous):

what formula do I use to solve: x^2 -50x +625 =0

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

x=25

OpenStudy (diyadiya):

Factorise

OpenStudy (anonymous):

\[x^2-2xy+y^2=(x-y)^2\] \[x^2-2(x)(25)+(25)^2=(x-25)^2\]

OpenStudy (anonymous):

(x-25)^2=0

OpenStudy (anonymous):

thanks!